# Quantum Hall Edges Beyond the Plasma Analogy

### v1

@@ Acknowledgments
for some fixed $\lambda$. By inspection, one can infer why the droplet must be finite: the integral in Eq. (49) grows logarithmically as $|\textbf{x}|\to\infty$ while the potential $W(\textbf{x})$ grows faster than that, so the equality cannot hold if $\Omega$ has non-compact support. Now, taking the Laplacian $\nabla^{2}=\partial_{x}^{2}+\partial_{y}^{2}$ of both sides of Eq. (49) gives
@@
Eq. (51) is useful to numerically simulate CGs of any given shape. For comparison with QH droplets, we restrict attention to CGs with uniform bulk density (50), i.e. CGs whose potential has constant Laplacian; see e.g. [51] for a discussion. Without loss of generality, we choose the droplet’s area to equal $\pi$, hence the normalized bulk density $\rho=1/\pi$. It is then convenient to use complex coordinates $z\equiv x^{1}+\mathrm{i}x^{2}$, $w\equiv y^{1}+\mathrm{i}y^{2}$ and turn the 2D integral (51) into a 1D integral thanks to Stokes’ theorem:
@@
Eq. (51) guarantees that this holds for $z\in\Omega$, but it also turns out to be a valid choice outside the droplet ($z\notin\Omega$). The key advantage of Eq. (52) is that it can be approximated numerically to high precision by a rectangle rule, so it provides a good approximation of the potential needed to simulate any given droplet shape. In the main text, Eq. (52) was used to produce the potential corresponding to flower droplets (19). The potential was then plugged into the energy (43) to eventually yield the Monte-Carlo data behind Fig. 5, similar to that in Ref. [58].
@@
$\langle\Phi(\textbf{y})\rangle^{\textbf{x}}\equiv\frac{\int\mathrm{d}\textbf{x}_{1}\ldots\mathrm{d}\textbf{x}_{N}\,\Phi(\textbf{y})\mathrm{e}^{-\beta E-\beta q\Phi(\textbf{x})}}{\int\mathrm{d}\textbf{x}_{1}\ldots\mathrm{d}\textbf{x}_{N}\,\mathrm{e}^{-\beta E-\beta q\Phi(\textbf{x})}}=\langle\Phi(\textbf{y})\rangle-q\beta\underbrace{\Big{(}\langle\Phi(\textbf{x})\Phi(\textbf{y})\rangle-\langle\Phi(\textbf{x})\rangle-\langle\Phi(\textbf{y})\rangle\Big{)}}_{\displaystyle\equiv\langle\Phi(\textbf{x})\Phi(\textbf{y})\rangle_{c}}+O(q^{2}).$ (55)
Here we expanded the mean up to first order in $q$, i.e. at the order of linear response, and we introduced the connected two-point function of $\Phi$. The result (55) was announced in the main text in Eq. (10). It says that correlations of the potential (53) determine the effects of small perturbations of the CG.
@@
    In this sense, the potential (53) behaves as a massless free boson in the bulk of a CG. We stress, however, that the relevant conformal structure will be outside the droplet, and ultimately has nothing to do with the conformal behavior of $\Phi(\textbf{x})$ as a free boson.
  • Now let x be outside the droplet while y lies in the bulk. Then the total potential $\langle\Phi(\textbf{y})\rangle^{\textbf{x}}-\langle\Phi(\textbf{y})\rangle-q\log|\textbf{x}-\textbf{y}|$ is that created by a perfect conductor with $N+1$ particles, and is therefore harmonic when seen as a function of x. Hence,
@@
Since the map (11) sends $\partial\Omega$ on a unit circle, one can label any point $\textbf{x}\in\partial\Omega$ by the angle $\alpha$ such that $G(\textbf{x})=\mathrm{e}^{\mathrm{i}\alpha}$. This is the “conformal” angle coordinate that governs the density-density correlations (12) on the edge of a CG. To derive these correlations, the key intuition stems again from electrostatics: as we saw, correlators of the field (53) are closely related to the electric potential generated at y (or $w$) by a point charge set at x (or $z$) in the presence of a conductor, the droplet $\Omega$. The resulting electric field is discontinuous on the boundary $\partial\Omega$ (see Fig. 2), resulting in a non-zero edge charge density $\sigma(\textbf{x})$ for $\textbf{x}\in\partial\Omega$. The main argument in [8] is that this phenomenon is responsible for the power-law decay of edge density correlations. Indeed, the edge charge density is
@@
where $\mathbf{n}_{\textbf{x}}$ is the outward normal unit vector at $\textbf{x}\in\partial\Omega$ and $\mathbf{E}_{\pm}(\textbf{x})\equiv\mathbf{E}(\textbf{x}\pm\epsilon\mathbf{n}_{\textbf{x}})$ are electric fields right outside and right inside the droplet. Since the electric field is $\mathbf{E}=-\nabla\Phi$, correlations of the charge density (61) can be related to those of $\Phi$ in Eqs. (56)–(58). The only non-zero contribution to the correlator $\langle\sigma(\textbf{x})\sigma(\textbf{y})\rangle_{c}$ turns out to stem from the Green’s function $K(\textbf{x},\textbf{y})$ in Eq. (58), and one finds
@@
A comment is in order regarding the conformal map (11) that plays a central role in the correlator (63). It turns out that the inverse map $G^{-1}:\mathbb{R}^{2}\backslash\mathbb{D}\to\mathbb{R}^{2}\backslash\Omega$ is often slightly easier to understand than the actual map itself, and can always be written as
@@
where the half-width is $e^{\lambda}$ and the half-height is $e^{-\lambda}$. Another is the square droplet (or more generally any polygonal droplet [50]), for which the inverse of (11) involves the Gauss hypergeometric function:
$G^{-1}(\zeta)=\frac{\Gamma(1/4)\,\zeta}{2\sqrt{2}\Gamma(3/4)}\,_{2}\!F_{1}(-1/2,-1/4;3/4,-1/\zeta^{4}).$ (66)
@@
$\textrm{Var}[\tilde{U}]_{\text{CG}}=-\frac{1}{2}\int_{\Omega}\mathrm{d}\textbf{r}\int\mathrm{d}\textbf{s}\left[\tilde{U}(\textbf{r}+\textbf{s}/2)-\tilde{U}(\textbf{r}-\textbf{s}/2)\right]^{2}\braket{\rho(\textbf{s})\rho(0)}_{c}\sim-\frac{1}{2}\int_{\Omega}\mathrm{d}\textbf{r}\int\mathrm{d}\textbf{s}\left[\textbf{s}\cdot\nabla\tilde{U}(\textbf{r})\right]^{2}\braket{\rho(\textbf{s})\rho(0)}_{c},$ (68)
@@
Here the edge correlator $\braket{\sigma(z)\sigma(w)}_{c}$ appears due to the double integral, in Eq. (67), along a direction perpendicular to the boundary $\partial\Omega$. Eq. (29) of the main text then follows from Eq. (71) upon using the edge correlator (63). While physically reasonable, the decoupling between bulk and edge contributions in Eq. (71) can also be checked by explicit computations at the free-fermion point $\beta=2$. It is ultimately justified by the fact that Eq. (29) can be proved with more advanced techniques; see [12].
@@
Part C: Symplectic versus conformal maps
@@
Here we address this puzzle within the class of edge-deformed potentials (18), with special focus on flower droplets. The corresponding action-angle coordinates are given by Eq. (17) with $R=0$ and a function $\theta(\varphi)$ given by Eq. (20). As we shall see, an extension from the canonical angle to a biholomorphism is possible for elliptic droplets ($k=2$), but impossible for droplets with $k\neq 2$ petals. Despite this, we will find that the local (not globally bijective) conformal extension does replicate the correlator of edge-deformed QH droplets.

### v2

@@ Acknowledgments
for some fixed $\lambda$. By inspection, one can infer why the droplet must be finite: the integral in Eq. (49) grows logarithmically as $|\textbf{x}|\to\infty$ while the potential $W(\textbf{x})$ grows faster than that, so the equality cannot hold if $\Omega$ has noncompact support. Now, taking the Laplacian $\nabla^{2}=\partial_{x}^{2}+\partial_{y}^{2}$ of both sides of Eq. (49) gives
@@
Eq. (51) is useful to numerically simulate CGs of any given shape. For comparison with QH droplets, we restrict attention to CGs with uniform bulk density (50), i.e. CGs whose potential has constant Laplacian; see e.g. [54] for a discussion. Without loss of generality, we choose the droplet’s area to equal $\pi$, hence the normalized bulk density $\rho=1/\pi$. It is then convenient to use complex coordinates $z\equiv x^{1}+\mathrm{i}x^{2}$, $w\equiv y^{1}+\mathrm{i}y^{2}$ and turn the surface integral (51) into a line integral thanks to Stokes’ theorem:
@@
Eq. (51) guarantees that this holds for $z\in\Omega$, but it also turns out to be a valid choice outside the droplet ($z\notin\Omega$). The key advantage of Eq. (52) is that it can be approximated numerically to high precision by a rectangle rule, so it provides a good approximation of the potential needed to simulate any given droplet shape. In the main text, Eq. (52) was used to produce the potential corresponding to flower droplets (19). The potential was then plugged into the energy (43) to eventually yield the Monte Carlo data behind Fig. 5, similar to that in [65].
@@
$\langle\Phi(\textbf{y})\rangle^{\textbf{x}}\equiv\frac{\int\mathrm{d}\textbf{x}_{1}\ldots\mathrm{d}\textbf{x}_{N}\,\Phi(\textbf{y})\mathrm{e}^{-\beta E-\beta q\Phi(\textbf{x})}}{\int\mathrm{d}\textbf{x}_{1}\ldots\mathrm{d}\textbf{x}_{N}\,\mathrm{e}^{-\beta E-\beta q\Phi(\textbf{x})}}=\langle\Phi(\textbf{y})\rangle-q\beta\underbrace{\Bigl{(}\langle\Phi(\textbf{x})\Phi(\textbf{y})\rangle-\langle\Phi(\textbf{x})\rangle\langle\Phi(\textbf{y})\rangle\Bigr{)}}_{\displaystyle\equiv\langle\Phi(\textbf{x})\Phi(\textbf{y})\rangle_{c}}+\,O(q^{2}).$ (55)
Here, we expanded the mean up to first order in $q$, i.e. at the order of linear response, and we introduced the connected two-point function of $\Phi$. The result (55) was announced in the main text in Eq. (10). It says that correlations of the potential (53) determine the effects of small perturbations of the CG.
@@
    In this sense, the potential (53) behaves as a massless free boson in the bulk of a CG. We stress, however, that the relevant conformal structure will be outside the droplet, and that it ultimately has nothing to do with the conformal behavior of $\Phi(\textbf{x})$ as a free boson.
  • Now let x be outside the droplet while y lies in the bulk. Then the total potential $\langle\Phi(\textbf{y})\rangle^{\textbf{x}}-\langle\Phi(\textbf{y})\rangle-q\log|\textbf{x}-\textbf{y}|$ is that created by a perfect conductor with $N+1$ particles, and it is therefore harmonic when seen as a function of x. Hence,
@@
Since the map (11) sends $\partial\Omega$ on a unit circle, one can label any point $\textbf{x}\in\partial\Omega$ by the angle $\alpha$ such that $G(\textbf{x})=\mathrm{e}^{\mathrm{i}\alpha}$. This is the conformal angle coordinate that governs the density-density correlations (12) on the edge of a CG. To derive these correlations, the key intuition stems again from electrostatics: as we saw, correlators of the field (53) are closely related to the electric potential generated at y (or $w$) by a point charge set at x (or $z$) in the presence of a conductor, the droplet $\Omega$. The resulting electric field is discontinuous on the boundary $\partial\Omega$ (see Fig. 2), resulting in a nonzero edge charge density $\sigma(\textbf{x})$ for $\textbf{x}\in\partial\Omega$. The main argument in [8] is that this phenomenon is responsible for the power-law decay of edge density correlations. Indeed, the edge charge density is
@@
where $\mathbf{n}_{\textbf{x}}$ is the outward normal unit vector at $\textbf{x}\in\partial\Omega$ and $\mathbf{E}_{\pm}(\textbf{x})\equiv\mathbf{E}(\textbf{x}\pm\epsilon\mathbf{n}_{\textbf{x}})$ are electric fields right outside and right inside the droplet. Since the electric field is $\mathbf{E}=-\nabla\Phi$, correlations of the charge density (61) can be related to those of $\Phi$ in Eqs. (56)–(58). The only nonzero contribution to the correlator $\langle\sigma(\textbf{x})\sigma(\textbf{y})\rangle_{c}$ turns out to stem from the Green’s function $K(\textbf{x},\textbf{y})$ in Eq. (58), and one finds
@@
A comment is in order regarding the conformal map (11) that plays a central role in the correlator (63). It turns out that the inverse map $G^{-1}:\mathbb{R}^{2}\backslash\mathbb{D}\to\mathbb{R}^{2}\backslash\Omega$ is often slightly easier to understand than the actual map itself, and that it can always be written as
@@
where the half-width is $e^{\lambda}$ and the half-height is $e^{-\lambda}$. Another is the square droplet (or more generally any polygonal droplet [53]), for which the inverse of (11) involves the Gaussian hypergeometric function:
$G^{-1}(\zeta)=\frac{\Gamma(1/4)\,\zeta}{2\sqrt{2}\Gamma(3/4)}\,_{2}F_{1}(-1/2,-1/4;3/4;-1/\zeta^{4}).$ (66)
@@
$\textrm{Var}[\tilde{U}]_{\text{CG}}=-\frac{1}{2}\int_{\Omega}\mathrm{d}\textbf{r}\int\mathrm{d}\textbf{s}\,\Bigl{[}\tilde{U}(\textbf{r}+\textbf{s}/2)-\tilde{U}(\textbf{r}-\textbf{s}/2)\Bigr{]}^{2}\!\braket{\rho(\textbf{s})\rho(0)}_{c}\sim-\frac{1}{2}\int_{\Omega}\mathrm{d}\textbf{r}\int\mathrm{d}\textbf{s}\,\Bigl{[}\textbf{s}\cdot\nabla\tilde{U}(\textbf{r})\Bigr{]}^{2}\!\braket{\rho(\textbf{s})\rho(0)}_{c},$ (68)
@@
Here, the edge correlator $\braket{\sigma(z)\sigma(w)}_{c}$ appears due to the double integral, in Eq. (67), along a direction perpendicular to the boundary $\partial\Omega$. Eq. (29) of the main text then follows from Eq. (71) upon using the edge correlator (63). While physically reasonable, the decoupling between bulk and edge contributions in Eq. (71) can also be checked by explicit computations at the free-fermion point $\beta=2$. It is ultimately justified by the fact that Eq. (29) can be proved with more advanced techniques; see [12].
@@
Part C: Symplectic vs conformal maps
@@
Here, we address this puzzle within the class of edge-deformed potentials (18), with special focus on flower droplets. The corresponding action-angle coordinates are given by Eq. (17) with $R=0$ and a function $\theta(\varphi)$ given by Eq. (20). As we shall see, an extension from the canonical angle to a biholomorphism is possible for elliptic droplets ($k=2$), but impossible for droplets with $k\neq 2$ petals. Despite this, we will find that the local (not globally bijective) conformal extension does replicate the correlator of edge-deformed QH droplets.